A 2uF capacitor is charged to 8V. It is then connected to a 12V supply and given additional charge

through a 1OMohms resistor. How long will it take to reach a voltage of 10V?


1
Expert's answer
2021-07-19T09:48:53-0400

A capacitor connected to a tension through a resistor satisfies the following equation :

"q/C+\dot q R=V" which is basically juste the expression of tension, using that the tension on the capacitor is "q\/C" and the tension on the resistance is "IR=\dot q R". Solving this equation yields :

"q/(RC)+\dot q = V/R"

"q=q_0 e^{-t\/(RC)}+CV"

By dividing both sides by "C" and using "U" for tension on the capacitor, we find

"U=U_0 e^{-t\/(RC)}+V"

We know that "U_{t=0}=8V". Therefore,

"8=U_0 +12", so "U_0 = -4V"

Now we will find the time when "U" reaches 10"V" :

"10=-4\cdot e^{-t/20}+12"

"e^{-t\/20}=1\/2"

"t=20\ln 2 \approx 13.86 \: s"